OpenStudy (mariebrittany07):

A plumber charges $55 per hour , plus a $9 service fee. If the plumber earns $225. How many hours did he work total ?

OpenStudy (anonymous):

3.92 that wat i got check it anyway

OpenStudy (anonymous):

@mariebrittany07 do you still want help with this one?

OpenStudy (mariebrittany07):

I added 55+9+255=319 that's what I had got for the answer and yes i still need help with this question

OpenStudy (anonymous):

To set it up, let's write an expression for how much the plumber charges: he charges $55 per hour, so for example, if the plumber worked 1 hour, he would charge 1 x $55... PLUS the $9 service fee. For 2 hours, he would charge (2 x 55) plus the $9 fee. You need to find out how many hours he worked... so we'll use that as our variable. H = # hours worked Then, the plumber charges 55 x H plus the $9 fee... 55H + 9 And the problem tells you his overall charge is $225. So, make an equation: 225 = 55H + 9

OpenStudy (mariebrittany07):

okay I did now what

OpenStudy (mariebrittany07):

are you still there jakeV8

OpenStudy (anonymous):

yes :)

OpenStudy (mariebrittany07):

so what do i do next do i flip it around

OpenStudy (mariebrittany07):

like this 9+55h=225

OpenStudy (mariebrittany07):

and can you please show me how you did it too

OpenStudy (anonymous):

You have an equation telling you: $225 is equal to $55 per hour multiplied by H hours, then plus $9 service fee. You need to solve for the number of hours, H. I'm sure you've done this sort of basic equation before: 225 = 55H + 9 Subtract 9 from both sides to get: 225-9 = 55H + 9 - 9 216 = 55H Then divide both sides by 55 to solve for H. 216/55 = (55/55)H = H H = 216/55

OpenStudy (anonymous):

you can put that in a calculator and get a decimal answer if you need to for the problem.

OpenStudy (anonymous):

yea she'll get 3.92 and a whole bunch of numbers

OpenStudy (anonymous):

do you see how to set up the equation though? Unless someone comes along and tells you the answer every time, you need to know how to set it up and solve for the missing variable.

OpenStudy (mariebrittany07):

okay

OpenStudy (mariebrittany07):

hey jake v8 i will be right back i have to eat

OpenStudy (anonymous):

enjoy :) I have to leave in a bit, but others will help you out :) good luck!

OpenStudy (mariebrittany07):

hey can someone help me PLEASE

OpenStudy (anonymous):

post a new question... I have to leave in about 5 minutes.

OpenStudy (mariebrittany07):

okay i still need help with this problem

OpenStudy (anonymous):

ok, that's fine... leave it open and "bump" it up on the stack... Good luck!

OpenStudy (mariebrittany07):

thanks a million

OpenStudy (mariebrittany07):

have a great day